OpenStudy (anonymous):

how can you tell if the points form a right triangle (2,4)(-1,6)(-3,1)??

OpenStudy (anonymous):

use point-slope form and write out the equations for the lines. then check to see whether any 2 lines are perpendicular with one another

OpenStudy (anonymous):

can you use the converse of the pythagorean theorem

OpenStudy (anonymous):

find the slope between each two points then if slope 1 * slope 2 =0 ===> the triangle is right triangle

OpenStudy (anonymous):

site has now told me at least 20 times that amistre gave me a medal for one problem. i mean thanks and all, but enough already!

OpenStudy (amistre64):

(2,4) (-1,6) (-3,1) -2-4 -2-4 -2-4 ------------------ 0,0 -3,2 -5,-3 aint that one (2,4) (-1,6) (-3,1) 1-6 1-6 1-6 ----------------- 3,-2 0,0 -2,-5 .... aint those (2,4) (-1,6) (-3,1) 3-1 3-1 3-1 ----------------- 5,3 2,5 0,0 aint those either

OpenStudy (anonymous):

find the distance between the points , call them a, b, c and check if \[a^2+b^2=c^2\]

OpenStudy (amistre64):

it keeps telling me you gave ME a medal ....

OpenStudy (amistre64):

ive refreshed, dumped cookies, logedin and out ...

OpenStudy (anonymous):

lol is that firebug underneath?

OpenStudy (amistre64):

chrome

OpenStudy (anonymous):

if 2 lines are perpendicular with one another, the product of their slopes is -1

OpenStudy (anonymous):

i mean underneath the web page. looks like firebug or something similar, where you can inspect the elements etc

OpenStudy (amistre64):

dunno, it a chrome feature; ctrl shift i

OpenStudy (anonymous):

ahhh

myininaya (myininaya):

so find the slopes of each of the lines and see if you get negative reciprocal when comparing two of the line segment slopes

myininaya (myininaya):

then it is a right triangle

OpenStudy (amistre64):

i compared vector components after putting each vertix at the origin

myininaya (myininaya):

but i think thats what amistre was saying above

myininaya (myininaya):

amistre you find the slope weird

OpenStudy (amistre64):

:) vectors are slopes .... so i just go for broke

myininaya (myininaya):

\[\frac{4-6}{2-(-1)}=\frac{-2}{3}, \frac{4-1}{2-(-3)}=\frac{3}{5}, \frac{6-1}{-1-(-3)}=\frac{5}{2}\] none of these are negative reciprocal of the other so no it is not a right triangle